# bibliography.bib

### v1

@@ Abstract
Assume that Theorem 3.4.1 holds for all $\textrm{GL}_{n^{\prime}},n^{\prime}<n,$ so that $\Phi_{n^{\prime}}^{k^{\prime}}$ are isomorphisms for all $k^{\prime}$, and the same holds also for $\textrm{GL}_{n}$, fixed $k$ and $\Phi_{n}^{k}$. Assume also that, for $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$, $\operatorname{Stab}_{W}\lambda\neq W$. Then
@@
Assume that $A\in D^{b}_{G}(G)$ satisfies $\mathfrak{hc}_{*}(A)\simeq\operatorname{Res}_{B}^{G}(A)$. Then there is a canonical natural isomorphisms of functors
@@
is supported on $\tilde{\mathcal{O}}_{1}$ (cf. (4) for notations). By Proposition 3.3.1,
@@
is an isomorphism. Proposition now follows from the base change isomorphism. ∎
We will also use the independence of $\operatorname{Res}^{G}_{P}:P_{G}(\mathcal{N}_{G})\to P_{G}(\mathcal{N}_{L})$ of the choice of the parabolic $P$ with Levi $L$, which is proved in [mirkovicCharacterSheavesReductive2004][Theorem 7.3], in the setting of sheaves on the nilpotent cone $\mathcal{N}_{\mathfrak{g}}$ of $\mathfrak{g}=\operatorname{Lie}G$. Since for a good prime $p$, $\mathcal{N}_{\mathfrak{g}}$ can be $G$-equivariantly identified with $\mathcal{N}_{G}$, see [humphreysConjugacyClassesSemisimple2011][Chapter 6.20] and references therein, same is true for equivariant perverse sheaves on $\mathcal{N}_{G}$. In this case we write $\operatorname{Res}^{G}_{P}=:\operatorname{Res}^{G}_{L}$.
@@
By Corollary 3.3.7, it is enough to show that
$\Pi_{\lambda}\operatorname{Res}_{Q}^{G}(\Phi_{n}^{k}\star\operatorname{id}_{E_{m}^{\lambda}})=\operatorname{Res}^{G}_{Q}(\Phi_{n}^{k}\star\operatorname{id}_{E_{m}^{\lambda}})$
is an isomorphism. This morphism, in turn, is, by definition, a direct summand of the morphism
@@
Choose a set of representatives $\Sigma$ of the double cosets $Q\backslash G/P$. For $x\in G$ write and a subgroup $H$ write ${}^{x}H=xHx^{-1},H^{x}=x^{-1}Hx$. We assume that $\Sigma$ is chosen so that $L$ and all $M^{x},x\in\Sigma$ have a common maximal torus. On the left hand side of (8), by Mackey formula, see [marsCharacterSheaves1989][Proposition 10.1.2], and Proposition 3.4.3, we get
$\Pi_{\lambda}\operatorname{Res}^{G}_{Q}\left(\operatorname{Ind}_{P}^{G}\operatorname{Res}^{G}_{P}(\mathcal{S}_{\wedge^{k}})\star^{0}E_{m}^{\lambda}\right)=\\ =\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}P}{}^{x}(\operatorname{Res}_{L\cap Q^{x}}^{L}\operatorname{Res}^{G}_{P}\mathcal{S}_{\wedge^{k}})\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=\\ =\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}{}^{x}(\operatorname{Res}_{M^{x}\cap L}^{G}\mathcal{S}_{\wedge^{k}})\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=\\ =\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}{}^{x}(\operatorname{Res}_{M^{x}\cap L}^{M^{x}}\operatorname{Res}^{G}_{M^{x}}\mathcal{S}_{\wedge^{k}})\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=\\ =\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}\operatorname{Res}_{M\cap{}^{x}L}^{M}{}^{x}(\operatorname{Res}^{G}_{M^{x}}\mathcal{S}_{\wedge^{k}})\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=\\ =\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}\operatorname{Res}_{M{}^{x}L}^{M}\operatorname{Res}^{G}_{M}\mathcal{S}_{\wedge^{k}}\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}.$
Note that since $P^{\lambda}(\mathcal{Y}_{Q})$ is a monoidal ideal containing $\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}$, we can omit $\Pi_{\lambda}$ from the notation.
@@
$\mathcal{S}_{\lambda}:=\Pi_{\lambda}\operatorname{Res}_{Q}^{G}(\mathcal{S}_{n}\star^{0}E_{m}^{\lambda})=\operatorname{Res}_{Q}^{G}(\mathcal{S}_{n})\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=\\ =\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}(\delta_{e}^{M\cap{}^{x}L})\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda},$
Here, for any Levi $N$, $\delta^{N}_{e}$ is a rank 1 skyscraper sheaf supported at the unit of $N$. We used the fact that $\mathcal{S}=\operatorname{Ind}_{P}^{G}\delta^{L}_{e}$ and Mackey formula.
By Corollary 3.3.7, on the right hand side of (8) we have
@@
Here we used the fact that $E_{\lambda}^{n}$ is an idempotent for $\star^{0}$, which follows from the corresponding fact for $G=T$.
@@
${\bigoplus\limits_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}P}\operatorname{Res}_{{}^{x}L\cap Q}^{M}\operatorname{Res}^{G}_{Q}\mathcal{S}_{\wedge^{k}}\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}}$${\mathcal{S}_{\lambda}\star^{0}\wedge^{k}\mathcal{S}_{\lambda}\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}}$$\scriptstyle{\tilde{\Phi}_{n,\lambda}^{k}}$
By the inductive assumption that $\Phi_{n^{\prime}}^{k}$ for all $n^{\prime}<n$ is an isomorphism, the restriction
$\Phi_{n,\lambda}^{k}=\Pi_{\lambda}\operatorname{Res}_{Q}^{G}(\Phi_{n}^{k}\star\operatorname{id}_{E_{m}^{\lambda}}):\operatorname{Res}^{G}_{Q}\mathcal{S}_{\wedge^{k+1}}\star^{0}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}\to\wedge^{k+1}\mathcal{S}_{\lambda}$
@@
Let $Z(H)$ stand for the center of an algebraic group $H$. For $c\in\mathcal{C}(Z(\textrm{GL}_{n}))$ let $e_{c}=\mathcal{E}_{1}^{c}[1]$ (in the notation of section 3.1.2 applied to $G=\mathbb{G}_{m}$). We will consider $e_{c}$ as an object in $P_{\textrm{GL}_{n}}(\textrm{GL}_{n})$ via the embedding $Z(\textrm{GL}_{n})\to\textrm{GL}_{n}$. As a direct consequence of the above and Corollary 3.3.4 we get
@@
Under the assumptions of the Proposition 3.4.2, perverse cohomology of the cone of $\operatorname{Id}_{e_{c}}\star\Phi_{n}^{k}:e_{c}\star\mathcal{S}_{\wedge^{k+1}}\to e_{c}\star\wedge^{k+1}\mathcal{S}_{n}$ are character sheaves with a central character fixed by $W$.
Note that it immediately follows that $\operatorname{Id}_{e_{c}}\star\Phi_{n}^{k}$ are isomorphisms for $k<n-1$: indeed, character sheaves on $\textrm{GL}_{n}$ have full support, while it is easy to see that the source and target of $\operatorname{Id}_{e_{c}}\star\Phi_{n}^{k}$ have closed support – product of $k+1<n$ matrices conjugate to a one in $U_{P}$ has an eigenvalue equal to 1 and lie in $\textrm{SL}_{n}$. It now follows from M1 that $\Phi_{n}^{k}$ is an isomorphism.
We now show how the cases $k\geq n-1$ follow from the following
@@
For any $c\in\mathcal{C}(Z(\textrm{GL}_{n}))$, $e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\mathcal{S}_{\wedge^{n-1}}$ does not contain character sheaves as direct summands.
Note that $\mathcal{S}_{n}\simeq\delta_{e}^{\textrm{SL}_{n}}\oplus\mathcal{C}_{n}$, where $\mathcal{C}_{n}$ is the intersection cohomology complex of the minimal non-unital unipotent orbit. It follows that
@@
It is easy to see that the functor $e_{c}\star^{0}-$ is monoidal and we have $\wedge^{k}\mathcal{S}_{c}\simeq e_{c}\star^{0}\wedge^{k}\mathcal{S}_{n}$. From Corollary 3.4.4 and discussion immediately below it, it follows that
@@
is a character sheaf, and so is $e_{c}\star^{0}\wedge^{n}\mathcal{C}_{n}$. On the other hand, $e_{c}\star^{0}\wedge^{n}\mathcal{C}_{n}\simeq e_{c}\star\wedge^{n}\mathcal{C}_{n}$ is the summand of $e_{c}\star^{0}\wedge^{n}\mathcal{S}_{n}$, which is in turn contained as a summand in
@@
Proposition 3.4.6.
Let $D$ be a regular semisimple matrix in $\textrm{SL}_{n}(\mathbf{k})$ with no eigenvalues equal to 1, let $P\subset\textrm{SL}_{n}$ be a parabolic subgroup fixing a line $l$. If a pair of matrices $M,N$, $M$ unipotent and $N\in U_{P}(\mathbf{k})$ with $NM=D$, exists, then $l$ contains a cyclic vector for $D$. If it exists, such a pair is unique for a given $l$.
@@
Note that if $l$ does not contain a cyclic vector, for any $N^{\prime}\in U_{P}$, $DN^{\prime}$ preserves an $(n-1)$-dimensional subspace and acts as a constant not equal to one on the quotient by this subspace, so can’t be unipotent.
Assume that $l$ contains a cyclic vector and that such a pair of matrices exists. Writing $N=1+A,M=1+B$, with $A$ and $B$ nilpotent, not that
$n=\operatorname{rk}(D-1)=\operatorname{rk}(AB+A+B)\leq\operatorname{rk}A+\operatorname{rk}B=1+\operatorname{rk}B.$
It follows that $M$ must be a regular unipotent element, such that $l$ is not in the image of $B$. Direct computation now shows that $N^{\prime}M$ with $N^{\prime}\in U_{P}(\mathbf{k})$ is unipotent if and only if $N^{\prime}=\operatorname{Id}$, hence the result. ∎
Proof of Proposition 3.4.5.
Let $\mathcal{N}$ stand for the unipotent variety of $\textrm{GL}_{n}$. The only summand of $e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\mathcal{S}_{\wedge^{n-1}}$ that can have full support is
$e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U].$
@@
$\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U]\simeq{}^{p}\mathcal{H}^{0}(m_{*}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathfrak{C}_{n}}[n^{2}+n-2]),$
Proposition 3.4.6 implies that the fiber of $m$ over the generic regular semisimple element in the intersection of the fixed torus $T$ with the image of $m$ is identified with $T_{adj}=T/Z(\textrm{GL}_{n})$. It follows that, if $e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U]$ contains a character sheaf as a direct summand, it must be a character sheaf with a generic stalk of dimension $1$. We are left to consider two cases, when restriction of this summand to $\textrm{SL}_{n}$ is a shifted constant sheaf, or a character sheaf corresponding to the Steinberg representation. Since both functors of *-convolution $\star$ and (non-equivariant) hypercohomology $\mathbb{H}^{0}$ are right t-exact, we have
@@
Finally, we show that the Steinberg character sheaf can not be a direct summand of $\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U]$, corresponding to $\wedge^{n+1}\mathcal{S}_{n}$. Recall the completed monodromic Hecke category $\mathcal{H}(\textrm{SL}_{n})$, the free pro-unipotent tilting object $\hat{\mathcal{T}}_{w_{0}}$, standard and costandard objects $\hat{\Delta}_{w},\hat{\nabla}_{w}$ in $\hat{D}$ from [bezrukavnikovKoszulDualityKacMoody2014]. We denote the composition
$D_{G}^{b}(G)\xrightarrow{\mathfrak{hc}}D_{G}^{b}(\mathcal{Y})\xrightarrow{\operatorname{For}_{T}}D_{G}^{b}(G/U\times G/U)\xrightarrow{~{}}D^{b}(U\backslash G/U)\to\hat{D}$
@@
Denote by $\pi:G/U\to G/B$ the canonical projection. Let $L_{w},w\in W$ stand for the pullback $\pi^{*}\operatorname{IC}_{w}$, where $IC_{w}$ is the intersection cohomology sheaf of the Bruhat $U$-orbit labeled by $w$. We collect some facts about the completed category and the functor $\mathfrak{hc}$ in the following
@@
    The functor $(-\star\hat{\mathcal{T}}_{w_{0}})$ of convolution with $\hat{\mathcal{T}}_{w_{0}}$ is t-exact with respect to the perverse t-structure.
@@
(a) follows from the monoidality of $\operatorname{For}_{T}$. (b) is proved in [bezrukavnikovCharacterDmodulesDrinfeld2012][Corollary 3.4] in the setting of $D$-modules and in [chenFormulaGeometricJacquet2017][Corollary 7.9] in $\ell$-adic setting. (c) is Lemma 4.4.6 with Lemma 4.4.11 (3) in [bezrukavnikovKoszulDualityKacMoody2014]. (d) is a direct consequence of (c). ∎
It is shown in [hhh], Corollary 6.2.7, that $\mathfrak{hc}^{\prime}(\mathcal{S}_{n})\star\hat{\Delta}_{1}$ has a filtration by the objects of the form $\mathfrak{J}_{w}\simeq\hat{\nabla}_{w}\star\hat{\nabla}_{w^{-1}}$ for $w\in W^{J}$. It follows from Proposition 2.1.1 (d) that
$\mathfrak{hc}^{\prime}(\mathcal{S}_{n})\star\hat{\mathcal{T}}_{w_{0}}\simeq V\otimes_{\overline{\mathbf{Q}}_{\ell}}\hat{\mathcal{T}}_{w_{0}},$
where $V$ is an $n$-dimensional $\overline{\mathbf{Q}}_{\ell}$-vector space. More generally,
$\mathfrak{hc}^{\prime}(\mathcal{S}_{n}^{\star k})\star\hat{\mathcal{T}}_{w_{0}}\simeq V^{\otimes k}\otimes_{\overline{\mathbf{Q}}_{\ell}}\mathcal{S}_{n}.$
@@
$\mathfrak{hc}^{\prime}(\mathcal{S}_{n}^{\star^{0}k})\simeq V^{\otimes k}\otimes_{\overline{\mathbf{Q}}_{\ell}}\mathcal{S}_{n},$
with $\Sigma_{k}$-action on the left hand side compatible with permutation action on $V^{\otimes k}$. The latter compatibility follows from the fact that $\mathfrak{hc}$ is a central functor, existence of the filtration by the objects ^$\hat{\nabla}_{w}\star\hat{\nabla}_{w^{-1}}$ of $\mathfrak{hc}(\mathcal{S}_{n})\star\hat{\nabla}_{1}$ and existence of non-zero morphisms ^$\hat{\nabla}_{1}\to\hat{\nabla}_{w}\star\hat{\nabla}_{w^{-1}}$.

### v2

@@ Abstract
Assume that Theorem 3.4.1 holds for all $\mathrm{GL}_{n^{\prime}},n^{\prime}<n,$ so that the $\Phi_{n^{\prime}}^{k^{\prime}}$ are isomorphisms for all $k^{\prime}$, and the same holds also for $\mathrm{GL}_{n}$, fixed $k$ and $\Phi_{n}^{k}$. Assume also that, for $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$, we have $\operatorname{Stab}_{W}\lambda\neq W$. Then
@@
Assume that $A\in D^{b}_{G}(G)$ satisfies $\mathfrak{hc}_{*}(A)\simeq\operatorname{Res}_{B}^{G}(A)$. Then there is a canonical natural isomorphism of functors
@@
is supported on $\tilde{\mathcal{O}}_{1}$ (see (4) for notations). By Proposition 3.3.1,
@@
is an isomorphism. The proposition now follows from the base change isomorphism. ∎
We will also use the independence of the functor $\operatorname{Res}^{G}_{P}:P_{G}(\mathcal{N}_{G})\to P_{G}(\mathcal{N}_{L})$ of the choice of a parabolic $P$ with the given Levi $L$, which is proved in [Mir04, Theorem 7.3] in the setting of sheaves on the nilpotent cone $\mathcal{N}_{\mathfrak{g}}$ of $\mathfrak{g}=\operatorname{Lie}G$. Since, for a good prime $p$, $\mathcal{N}_{\mathfrak{g}}$ can be $G$-equivariantly identified with $\mathcal{N}_{G}$ (see [Hum11, Chapter 6.20] and references therein), the same is true for equivariant perverse sheaves on $\mathcal{N}_{G}$. In this case we write $\operatorname{Res}^{G}_{P}=:\operatorname{Res}^{G}_{L}$.
@@
By Corollary 3.3.7, it is enough to show that $\Pi_{\lambda}\operatorname{Res}_{Q}^{G}(\Phi_{n}^{k}\star\operatorname{id}_{E_{m}^{\lambda}})$ is an isomorphism for all $m\geq 1$. This morphism is, by definition, a direct summand of the morphism
@@
[…]
$\Pi_{\lambda}\operatorname{Res}^{G}_{Q}$ $\left(\operatorname{Ind}_{P}^{G}\operatorname{Res}^{G}_{P}(\mathcal{S}_{\wedge^{k}})\star^{0}E_{m}^{\lambda}\right)=$
$=\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}P}{}^{x}(\operatorname{Res}_{L\cap Q^{x}}^{L}\operatorname{Res}^{G}_{P}\mathcal{S}_{\wedge^{k}})\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=$
$=\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}{}^{x}(\operatorname{Res}_{M^{x}\cap L}^{G}\mathcal{S}_{\wedge^{k}})\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=$
$=\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}{}^{x}(\operatorname{Res}_{M^{x}\cap L}^{M^{x}}\operatorname{Res}^{G}_{M^{x}}\mathcal{S}_{\wedge^{k}})\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=$
$=\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}\operatorname{Res}_{M\cap{}^{x}L}^{M}{}^{x}(\operatorname{Res}^{G}_{M^{x}}\mathcal{S}_{\wedge^{k}})\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}=$
$=\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}\operatorname{Res}_{M{}^{x}L}^{M}\operatorname{Res}^{G}_{M}\mathcal{S}_{\wedge^{k}}\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}.$
Note that since $P^{\lambda}(\mathcal{Y}_{Q})$ is a monoidal ideal containing $\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}$, we can omit $\Pi_{\lambda}$ on the first $\star^{0}$-factor from the notation.
@@
$\mathcal{S}_{\lambda}:=\Pi_{\lambda}\operatorname{Res}_{Q}^{G}(\mathcal{S}_{n}\star^{0}E_{m}^{\lambda})$ $=\operatorname{Res}_{Q}^{G}(\mathcal{S}_{n})\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}$
$=\bigoplus_{x\in\Sigma}\operatorname{Ind}^{M}_{M\cap{}^{x}L}(\delta_{e}^{M\cap{}^{x}L})\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}.$
Here, for any Levi $N$, $\delta^{N}_{e}$ is the rank 1 skyscraper sheaf supported at the unit of $N$. Above we used the fact that $\mathcal{S}=\operatorname{Ind}_{P}^{G}\delta^{L}_{e}$ and the Mackey formula.
By Corollary 3.3.7, on the right-hand side of (8) we have
@@
Here we used the fact that $E^{\lambda}_{m}$ is an idempotent for $\star^{0}$, which follows from the corresponding fact for $G=T$.
@@
${\bigoplus\limits_{x\in\Sigma}}\operatorname{Ind}^{M}_{M\cap{}^{x}P}\operatorname{Res}_{{}^{x}L\cap Q}^{M}\operatorname{Res}^{G}_{Q}\mathcal{S}_{\wedge^{k}}\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}$$\mathcal{S}_{\lambda}\star^{0}\wedge^{k}\mathcal{S}_{\lambda}\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}$$\scriptstyle{\tilde{\Phi}_{n,\lambda}^{k}}$
By the inductive assumption that $\Phi_{n^{\prime}}^{k}$ is an isomorphism for all $n^{\prime}<n$, the restriction
$\Phi_{n,\lambda}^{k}=\Pi_{\lambda}\operatorname{Res}_{Q}^{G}(\Phi_{n}^{k}\star\operatorname{id}_{E_{m}^{\lambda}}):\operatorname{Res}^{G}_{Q}\mathcal{S}_{\wedge^{k+1}}\star^{0}\Pi_{\lambda}\operatorname{Res}^{G}_{Q}E_{m}^{\lambda}\to\wedge^{k+1}\mathcal{S}_{\lambda}$
@@
Let $Z(H)$ stand for the center of an algebraic group $H$. For $c\in\mathcal{C}(Z(\textrm{GL}_{n}))(\overline{\mathbf{Q}}_{\ell})$, let $e_{c}=\mathcal{E}_{1}^{c}[1]$ (in the notation of Section 3.1.2 applied to $G=\mathbb{G}_{m}$). We will consider $e_{c}$ as an object in $P_{\textrm{GL}_{n}}(\textrm{GL}_{n})$ via the closed embedding $Z(\textrm{GL}_{n})\to\textrm{GL}_{n}$. As a direct consequence of the above and Corollary 3.3.4 we get
@@
Under the assumptions of Proposition 3.4.2, perverse cohomology sheaves of the cone of the map
$\operatorname{Id}_{e_{c}}\star\,\Phi_{n}^{k}:e_{c}\star\mathcal{S}_{\wedge^{k+1}}\to e_{c}\star\wedge^{k+1}\mathcal{S}_{n}$
are character sheaves with a central character fixed by $W$.
Note that it immediately follows that the morphisms $\operatorname{Id}_{e_{c}}\star\,\Phi_{n}^{k}$ are isomorphisms for k<n1k<n-1: indeed, character sheaves on $\textrm{GL}_{n}$ have full support, while it is easy to see that the source and target of $\operatorname{Id}_{e_{c}}\star\,\Phi_{n}^{k}$ have closed support, since a product of $k+1<n$ matrices conjugate to a one in $U_{P}$ has an eigenvalue equal to 1 and lies in $\textrm{SL}_{n}$. It now follows from M1 that $\Phi_{n}^{k}$ is an isomorphism.
We now deduce the case $k\geq n-1$ from the following
@@
For any $c\in\mathcal{C}(Z(\mathrm{GL}_{n}))$, $e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\mathcal{S}_{\wedge^{n-1}}$ does not contain character sheaves as direct summands.
Note that $\mathcal{S}_{n}\simeq\delta_{e}\oplus\mathcal{C}_{n}$, where $\mathcal{C}_{n}$ is the intersection cohomology complex of the minimal unipotent orbit. It follows that
@@
It is easy to see that the functor $e_{c}\star^{0}$- is monoidal and we have $\wedge^{k}(e_{c}\star^{0}\mathcal{S}_{n})\simeq e_{c}\star^{0}\wedge^{k}\mathcal{S}_{n}$. From Corollary 3.4.4 and the discussion immediately below it, it follows that
@@
is a character sheaf, and so is $e_{c}\star^{0}\wedge^{n}\mathcal{C}_{n}$. On the other hand, we have that the object $e_{c}\star^{0}\wedge^{n}\mathcal{C}_{n}\simeq e_{c}\star\wedge^{n}\mathcal{C}_{n}$ is a summand of $e_{c}\star^{0}\wedge^{n}\mathcal{S}_{n}$, which is in turn contained as a summand in
@@
Lemma 3.4.6.
Let $D$ be a regular semisimple matrix in $\mathrm{SL}_{n}(\mathbf{k})$ with no eigenvalues equal to 1. Let $P\subset\mathrm{SL}_{n}$ be the parabolic subgroup fixing a line $l$. If a pair of matrices $M,N,$ with $M$ unipotent, $N\in U_{P}(\mathbf{k})$, and $NM=D$, exists, then $l$ contains a cyclic vector for $D$. If it exists, such a pair is unique for a given $l$.
@@
Note that if $l$ does not contain a cyclic vector, then for any $N^{\prime}\in U_{P}$, $N^{\prime}D$ preserves an (n1)(n-1)-dimensional subspace and acts as a constant not equal to 1 on the quotient by this subspace, so can’t be unipotent.
Assume that $l$ contains a cyclic vector and that such a pair of matrices $M,N$ exists. Writing $N=\operatorname{id}+A$ and $M=\operatorname{id}+B$, with $A$ and $B$ nilpotent, note that
$n=\operatorname{rk}(D-\operatorname{id})=\operatorname{rk}(AB+A+B)\leq\operatorname{rk}A+\operatorname{rk}B=1+\operatorname{rk}B.$
It follows that $M$ must be a regular unipotent element, such that $l$ is not in the image of $B$. Direct computation now shows that $N^{\prime}M$ with $N^{\prime}\in U_{P}(\mathbf{k})$ is unipotent if and only if $N^{\prime}=\operatorname{id}$. If $M_{1},N_{1}$ is another such pair satisfying $N_{1}M_{1}=D,$ we have that $M_{1}=(N_{1}^{-1}N)M$ is unipotent, so that $N_{1}=N$ and $M_{1}=M$, as desired. ∎
Since $e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\mathcal{S}_{\wedge^{n-1}}$ is the direct sum of perverse sheaves with proper closed support and
$e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U],$
any direct summand of the former sheaf with full support is also a direct summand of the latter.
@@
$\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U]\simeq{}^{p}\mathcal{H}^{0}(m_{*}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathfrak{C}_{n}}[n^{2}+n-2]).$
Lemma 3.4.6 implies that the fiber of $m$ over the generic regular semisimple element in the intersection of the fixed torus $T$ with the image of $m$ is identified with $T_{adj}=T/Z(\textrm{GL}_{n})$. It follows that, if $e_{c}\star^{0}\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U]$ contains a character sheaf as a direct summand, it must be a character sheaf with a generic stalk of dimension $1$. We are left to consider two cases, when restriction of this summand to $\textrm{SL}_{n}$ is a shifted constant sheaf, or a character sheaf corresponding to the Steinberg representation. Since the functors of $\ast$-convolution and (non-equivariant) hypercohomology $\mathbb{H}^{0}$ are both right $t$-exact, we have
@@
Finally, we show that the Steinberg character sheaf can not be a direct summand of $\mathcal{S}_{n}\star^{0}\underline{\overline{\mathbf{Q}}_{\ell}}_{\mathcal{N}}[2\dim U]$, corresponding to $\wedge^{n+1}\mathcal{S}_{n}$. We recall some notations from Section 2. There we defined the completed monodromic Hecke category $\mathcal{H}(\textrm{SL}_{n})$, the free-monodromic tilting object $\hat{\mathcal{T}}_{w_{0}}$, standard and costandard objects $\hat{\Delta}_{w},\hat{\nabla}_{w}$ in $\mathcal{H}(\textrm{SL}_{n})$ from [BY13]. We denote the composition
[…]
@@
Denote by $\pi:G/U\to G/B$ the canonical projection. Let $L_{w}$ for $w\in W$ stand for the pullback $\pi^{*}\operatorname{IC}_{w}$, where $\operatorname{IC}_{w}$ is the intersection cohomology sheaf of the Bruhat $U$-orbit labeled by $w$. We collect some facts about the completed category and the functor $\mathfrak{hc}$ in the following
@@
    The functor $(-)\star\hat{\mathcal{T}}_{w_{0}}$ of convolution with $\hat{\mathcal{T}}_{w_{0}}$ is $t$-exact with respect to the perverse $t$-structure.
@@
(a) follows from the monoidality of $\operatorname{For}_{T}$. (b) is proved in [BFO12, Corollary 3.4] in the setting of $D$-modules and in [CY17, Corollary 7.9] in the $\ell$-adic setting. (c) is [BY13, Lemma 4.4.6, Lemma 4.4.11 (3)]. (d) is a direct consequence of (c). ∎
It is shown in [BT22], Corollary 6.2.7, that $\mathfrak{hc}^{\prime}(\mathcal{S}_{n})\star\hat{\Delta}_{e}$ has a filtration by the objects of the form $\mathfrak{J}_{w}\vcentcolon=\hat{\nabla}_{w}\star\hat{\nabla}_{w^{-1}}$ for $w\in W^{J}$. It follows from Proposition 2.1.1 (d) that
$\mathfrak{hc}^{\prime}(\mathcal{S}_{n})\star\hat{\mathcal{T}}_{w_{0}}\simeq{\mathbf{V}}\otimes_{\overline{\mathbf{Q}}_{\ell}}\hat{\mathcal{T}}_{w_{0}},$
where $\bf{V}$ is an $n$-dimensional $\overline{\mathbf{Q}}_{\ell}$-vector space that can be identified with $\mathbf{K}^{n}$ using the isomorphism $\mathfrak{J}_{w}\star\hat{\mathcal{T}}_{w_{0}}\simeq\hat{\mathcal{T}}_{w_{0}}$. More generally,
$\mathfrak{hc}^{\prime}(\mathcal{S}_{n}^{\star k})\star\hat{\mathcal{T}}_{w_{0}}\simeq\mathbf{V}^{\otimes k}\otimes_{\overline{\mathbf{Q}}_{\ell}}\mathcal{S}_{n}.$
@@
$\mathfrak{hc}^{\prime}(\mathcal{S}_{n}^{\star^{0}k})\simeq\mathbf{V}^{\otimes k}\otimes_{\overline{\mathbf{Q}}_{\ell}}\mathcal{S}_{n},$
with $S_{k}$-action on the left-hand side compatible with permutation action on $\mathbf{V}^{\otimes k}$.